OpenStudy (anonymous):

Write a rule for the linear function in the graph (4,3) and (3,-1) A. y = -4x + 13 B. y = -4x – 13 C. y = 4x – 13 D. y = 1/4x – 13

jimthompson5910 (jim_thompson5910):

first find the slope of the line through those two points

jimthompson5910 (jim_thompson5910):

Use the slope formula m = (y2-y1)/(x2-x1) m = (-1-3)/(3-4) m = (-4)/(-1) m = 4

jimthompson5910 (jim_thompson5910):

now find the y-intercept y = mx + b y = 4x + b 3 = 4(4) + b solve for b

OpenStudy (dan815):

hey

OpenStudy (dan815):

lemme show you how to get line equation, first u need slope of line

OpenStudy (dan815):

draw these 2 points of the graph if that helps you see it

OpenStudy (dan815):